OpenStudy (anonymous):

PLEASE HELP: Find the values of the trigonometric functions of from the given information. sin t =-3/5 t is in quadrant 3 cos t= tan= csc= sec= cot=

OpenStudy (anonymous):



OpenStudy (anonymous):

other side is 4, so take the appropriate ratios, remembering that you are in quadrant 3 so sine will be negative, cosine will be negative, tangent will be positive etc

OpenStudy (anonymous):

this is what I got but I don't know which should be negative or positive cos=4/5 tan=-3/4 csc=-5/3 sec=4/5 cot=-4/3

OpenStudy (anonymous):

can you check thos for me?

OpenStudy (anonymous):

quadrant 3 cosine will be negative

OpenStudy (anonymous):

tangent will be postive

OpenStudy (anonymous):

cos= -4/5 tan= 3/4 csc= -5/3 sec= -4/5 cot= 4/3

OpenStudy (anonymous):

okay thanks for the clarification, can you help me on onemore?